# Pixel-level and Semantic-level Adjustable Super-resolution: A Dual-LoRA Approach

### v1

@@ Abstract
Diffusion prior-based methods have shown impressive results in real-world image super-resolution (SR). However, most existing methods entangle pixel-level and semantic-level SR objectives in the training process, struggling to balance pixel-wise fidelity and perceptual quality. Meanwhile, users have varying preferences on SR results, thus it is demanded to develop an adjustable SR model that can be tailored to different fidelity-perception preferences during inference without re-training. We present Pixel-level and Semantic-level Adjustable SR (PiSA-SR), which learns two LoRA modules upon the pre-trained stable-diffusion (SD) model to achieve improved and adjustable SR results. We first formulate the SD-based SR problem as learning the residual between the low-quality input and the high-quality output, then show that the learning objective can be decoupled into two distinct LoRA weight spaces: one is characterized by the $\ell_{2}$-loss for pixel-level regression, and another is characterized by the LPIPS and classifier score distillation losses to extract semantic information from pre-trained classification and SD models. In its default setting, PiSA-SR can be performed in a single diffusion step, achieving leading real-world SR results in both quality and efficiency. By introducing two adjustable guidance scales on the two LoRA modules to control the strengths of pixel-wise fidelity and semantic-level details during inference, PiSA-SR can offer flexible SR results according to user preference without re-training. Codes and models can be found at https://github.com/csslc/PiSA-SR.
@@ 1 Introduction
Figure 1: Visual illustration of our pixel- and semantic-level adjustable method for real-world super-resolution. By increasing the guidance scale $\lambda_{pix}$ on the pixel-level LoRA module, the image degradations such as noise and compression artifacts can be gradually removed; however, a too-strong $\lambda_{pix}$ will make the SR image over-smoothed. By increasing the guidance scale $\lambda_{sem}$ on the semantic-level LoRA module, the SR images will have more semantic details; nonetheless, a too-high $\lambda_{sem}$ will generate visual artifacts. Please zoom in for a better view.
@@
Recently developed large-scale stable diffusion (SD) models [35], which are pre-trained on text-to-image (T2I) tasks, have demonstrated impressive image generation capabilities with strong semantic understanding and have been successfully applied to many downstream tasks [64, 33], including single-image SR [43, 28, 57, 54, 53]. While SD-based SR approaches have exhibited perceptually more realistic SR results than GAN-based methods, they commonly entangle the pixel-level fidelity and semantic-level enhancement objectives [43, 54, 53], which can be contradictory in optimization [3], in the diffusion process. As a result, they struggle to balance the pixel-wise fidelity and semantic-level perception in the final SR output. While some methods [28, 57, 40] design to perform pixel-level restoration before semantic-level enhancement, i.e., in a sequential manner, the performance of the latter stage depends heavily on the accuracy of the earlier stage.
@@
Besides the trade-off of pixel-level fidelity and semantic-level details, users in practice often have different preferences on the SR results: some prioritize content fidelity over detail generation, while others favor semantically rich details over pixel-wise fidelity. This diversity in user preferences underscores the importance and demands for a more flexible SR approach to accommodate individual tastes during inference. Although some previous methods [9, 32] have explored interactive SR methods to control the restoration strength, they are primarily limited to degradation-level adjustments (e.g., noise and blur), resulting in either sharper or smoother outputs. Several multi-step SD-based SR methods [54, 57, 28, 59] employ guidance-based strategies [12] at each sampling step to achieve varying levels of semantic richness, but these methods often struggle with precise control and efficiency.
@@
In this paper, we present Pixel-level and Semantic-level Adjustable Super-Resolution, namely PiSA-SR, which separates pixel-level and semantic-level enhancement into two distinct Low-Rank Adapter (LoRA) [14] weight spaces with the pre-trained SD model, providing an effective SR model for various user preferences. We begin by formulating the SD-based SR problem to learn the residual between LQ and HQ latent features [10, 62, 67]. With this formulation, we can not only accelerate the convergence of model training but also introduce scale factors on the model outputs, enabling flexible adjustment on the SR results during the inference stage without re-training. Then, we fine-tune the pre-trained SD model by introducing two separate LoRA modules dedicated to pixel-level regression and semantic-level enhancement. In specific, the $\ell_{2}$ loss is employed for the LoRA module of pixel-level regression, while the LPIPS [65] and classifier score distillation (CSD) [60] losses are employed for another LoRA module, leveraging the semantic priors encoded in the pre-trained VGG classification model [37] and SD image generation model [35].
A decoupled training approach is proposed to train pixel- and semantic-level LoRAs, effectively improving semantic information while maintaining pixel-level fidelity.
Our experiments demonstrate that PiSA-SR not only achieves superior SR performance to existing SD-based models but also offers an effective way for users to adjust the SR styles based on their tastes. An example is shown in Fig. 1.
@@
The horizontal and vertical axes indicate the semantic-level and pixel-level enhancement scales, respectively. Increasing the pixel-level enhancement scale effectively removes noise and compression artifacts, but a too-strong scale over-smoothes image details. Conversely, increasing the semantic-level enhancement scale enriches the image details, making the image perceptually realistic. Nonetheless, a too-high semantic enhancement scale will introduce over-enhanced visual artifacts. PiSA-SR provides the flexibility to adjust both pixel-level and semantic-level enhancement scales based on user preferences.
@@ 2 Related Work
Image super-resolution. Early deep learning-based SR methods [5, 8, 20, 25] aim to improve image fidelity metrics such as PSNR and SSIM [49]. By using bicubic downsampling to generate LQ-HQ image training pairs, various approaches have been developed to enhance the SR performance, such as dense [41], residual [67], recursive connections [20], non-local strategies [44], and attention mechanisms [25, 66, 6]. Though these advancements have significantly improved the SR performance, the real-world LQ images with complex degradations are much more challenging to enhance. While researchers have proposed to collect real-world LQ-HQ paired datasets [4, 52] to train SR models, simulating realistic LQ-HQ pairs is a more economic way [63, 47]. BSRGAN [63] randomly shuffles some basic degradation operators to synthesize LQ-HQ pairs, while RealESRGAN [47] implements high-order degradation modeling. Compared with pixel-level losses, the perceptual loss [15] and GAN loss [22, 45] are more effective in improving image perceptual quality. However, the adversarial training also leads to unnatural visual artifacts in the SR results. Many of the later works aim to reduce GAN-generated artifacts from the perspectives of frequency division [23], network optimization [68], the weighting of different losses [46], and image local statistics [27].
@@
Diffusion model based super-resolution. Recent studies have explored the use of diffusion models (DMs) [13, 35] for SR. Early approaches adjust pre-trained DMs’ reverse transitions [18] using gradient descent, offering training-free solutions by assuming predefined image degradations. ResShift [61] trains a DM from scratch on paired LQ-HQ data. Recently, SD models pre-trained on T2I tasks have been widely used for SR tasks due to their strong image priors [43, 28, 57, 54, 40, 53, 48, 55, 30, 2, 34]. StableSR [43] introduces a trainable encoder and takes LQ image as a condition in SD. DiffBIR [28] first applies a restoration module to mitigate degradations, then utilizes SD to enhance details. PASD [57] employs an encoder for degradation removal and introduces a pixel-aware cross-attention module to incorporate low-level and high-level image features into the SD process. SeeSR [54] enhances semantic robustness through degradation-aware tag-style prompts to guide diffusion. Considering that the multi-step diffusion process will increase the computational cost and the risk of synthesizing unfaithful contents, researchers started to develop fewer-step DM-based SR models. SinSR [48] applies consistency-preserving distillation to shorten the diffusion process of ResShift. OSEDiff [53] employs the LQ image as direct input, eliminating the random noise sampling. It adopts the VSD loss [51, 58] to distill the generative capacity from the multi-step diffusion process, offering a one-step DM-based SR solution.
However, all the above-mentioned methods face a contradiction between pixel-level and semantic-level enhancement and most of them lack the flexibility to adjust the SR style for diverse user preferences.
@@ 3.1 Model Formulation
Learning the residual between the LQ and GT features has been successfully applied in deep learning-based restoration methods [62, 67]. However, DMs typically perform multi-step iterations, making conventional residual learning difficult to apply. The recently developed OSEDiff [53] method employs a single diffusion step to achieve end-to-end SR training, making it feasible to adopt the residual learning strategy. Here, we formulate the SR problem as learning the residual between $z_{L}$ and $z_{H}$, as illustrated in Fig. 2 (c). Because DMs are trained to denoise in the inverse diffusion process, we use subtraction in the global residual connection. Such a residual learning formulation helps the model focus on learning the desired high-frequency information from the HQ latent features, simplifying the task of extracting less relevant information from the LQ latent. It can also accelerate the convergence of the model training process [10]. Furthermore, during inference, we can introduce a scaling factor $\lambda$ to adjust the residual $\epsilon_{\theta}$ added to the LQ latent, which can be expressed as:
@@
Users can adjust the output by using a smaller $\lambda$ (i.e., preserving more original contents) or a higher $\lambda$ (i.e., enhancing details more aggressively) based on their preferences.
@@ 3.2 Dual-LoRA Training
Previous SD-based SR methods mostly entangle the pixel-level and semantic-level enhancement in the training process [43, 53], making the balance of content fidelity and perpetual quality difficult. Some works sequentially perform pixel-level and semantic-level enhancement using two separate networks [28]. However, the errors generated in the earlier stage may propagate to the subsequent stage, limiting the overall performance. Additionally, using two separate networks increases the computational burden and memory requirements, making the approach less efficient. It has been successfully applied in customized T2I tasks by using distinct LoRA modules for various generation styles [36, 16, 39]. Inspired by this, we propose a decoupled training approach by utilizing two LoRA modules under the pre-trained SD model, targeting pixel-level and semantic-level enhancement for the SR task respectively. Our approach, namely Dual-LoRA, introduces only a small number of additional parameters during training. These LoRA parameters can also be merged into the pre-trained SD model during inference.
@@
Figure 4: The outputs of the model trained with pixel-wise and semantic-level losses for a given LQ image.
@@ 3.3 Pixel-level and Semantic-level Losses
where $c$ is the text prompt, $t$ is the current timestep, and $\sigma_{t}=\sqrt{1-{{\bar{\alpha}}_{t}}}$. The SD model, parameterized by $\theta$, model score functions of $\nabla_{x_{t}}\log p_{\theta}(x_{t}\mid c)$ and $\nabla_{x_{t}}\log p_{\theta}(x_{t})$ using noise predictions [38]. By taking the expectation of the gradient in Eq. 4 over all possible $t$, we have the CSD [12] loss, which is originally used in 3D generation tasks to optimize the posterior probability of rendered images aligning their semantic content with text prompts.
@@
Note that the VSD loss [51, 58], which also aligns the distributions of enhanced and natural images in the latent space, has been verified effective for SR by OSEDiff [53]. The gradient of VSD loss can be written as:
@@ 3.4 The Inference Process of PiSA-SR
Fig. 3(b) illustrates the inference process of PiSA-SR. In the default inference setting (represented by the green dash-dot line), only the PiSA-LoRA, which merges the pixel-level and semantic-level LoRA modules, is used to process the input together with the pre-trained SD model, achieving state-of-the-art SR performance in just one-step diffusion (see our experiments in Section 4.3). To achieve flexible SR with diverse user preferences, inspired by CFG [12], we introduce a pair of pixel and semantic guidance factors, denoted by $\lambda_{pix}$ and $\lambda_{sem}$, to control the SR results as follows:
@@
In Eq. (8), $\epsilon_{\theta_{pix}}(z_{L})$ is the output with only pixel-level LoRA, while $\epsilon_{\theta_{P\!i\!S\!A}}(z_{L})$ is the output with both pixel and semantic level enhancement. The difference between them, i.e., $\epsilon_{\theta_{P\!i\!S\!A}}(z_{L})-\epsilon_{\theta_{pix}}(z_{L})$, can well represent the enhanced semantic details (please refer to the visualization in Fig. 4). Therefore, as illustrated in Fig. 3(b) (represented by the black dashed line), by adjusting the values of $\lambda_{pix}$ and $\lambda_{sem}$, we can control the contributions of pixel-level LoRA and semantic-level LoRA, and consequently deliver the SR results with different styles, as illustrated in Fig. 1.
@@ 4.1 Experimental Settings
Training settings. We train PiSA-SR upon SD 2.1-base for the ×$\times 4$ SR task. Following SeeSR [54] and OSEDiff [53], we employ LSDIR [24] and the first 10K images from FFHQ [17] as the training data. We generate paired training data using the degradation pipeline from RealESRGAN [47]. The batch size is 16 and the training patch size is 512×512. The Adam optimizer [21] is used with a learning rate of 5e-5. The pixel-level and semantic-level LoRA modules undergo 4K and 8.5K iterations of training.
@@
Test datasets and evaluation metrics. Following previous arts [43, 54, 53], we evaluate the competing methods using synthetic and real-world test data. The synthetic dataset comprises 3000 images cropped from DIV2K [1] with size $512\times 512$, degraded using RealESRGAN degradation pipeline [47]. The real-world data are center-cropped from RealSR [4] and DrealSR [52] datasets with size $128\times 128$ for LQ images and $512\times 512$ for HQ images. PSNR and SSIM [49], computed on the Y channel in the YCbCr space, are used to measure the fidelity of SR results; LPIPS [65] and DISTS [7], computed in the RGB space, are used to assess the perceptual quality of SR results; FID [11] evaluates the distance of distributions between GT and restored images; NIQE [31], CLIPIQA [42], MUSIQ [19] and MANIQA [56] evaluate the image quality without reference images.
In the following, we first demonstrate the capability of pixel- and semantic-level adjustable SR of PiSA-SR, then compare it (with default setting) with competing methods, followed by complexity analysis and ablation studies.

### v2

@@ Abstract
Diffusion prior-based methods have shown impressive results in real-world image super-resolution (SR). However, most existing methods entangle pixel-level and semantic-level SR objectives in the training process, struggling to balance pixel-wise fidelity and perceptual quality. Meanwhile, users have varying preferences on SR results, thus it is demanded to develop an adjustable SR model that can be tailored to different fidelity-perception preferences during inference without re-training. We present Pixel-level and Semantic-level Adjustable SR (PiSA-SR), which learns two LoRA modules upon the pre-trained stable-diffusion (SD) model to achieve improved and adjustable SR results. We first formulate the SD-based SR problem as learning the residual between the low-quality input and the high-quality output, then show that the learning objective can be decoupled into two distinct LoRA weight spaces: one is characterized by the $\ell_{2}$-loss for pixel-level regression, and another is characterized by the LPIPS and classifier score distillation losses to extract semantic information from pre-trained classification and SD models. In its default setting, PiSA-SR can be performed in a single diffusion step, achieving leading real-world SR results in both quality and efficiency. By introducing two adjustable guidance scales on the two LoRA modules to control the strengths of pixel-wise fidelity and semantic-level details during inference, PiSA-SR can offer flexible SR results according to user preference without re-training. The source code of our method can be found at https://github.com/csslc/PiSA-SR.
@@ 1 Introduction
Figure 1: Visual illustration of our pixel- and semantic-level adjustable method for real-world SR. By increasing the pixel-level guidance scale $\lambda_{pix}$, the image degradations such as noise and compression artifacts can be gradually removed; however, a too-strong $\lambda_{pix}$ will make the SR image over-smoothed. By increasing the semantic-level guidance scale $\lambda_{sem}$, the SR images will have more semantic details; nonetheless, a too-high $\lambda_{sem}$ will generate visual artifacts. Please zoom in for a better view.
@@
Recently developed large-scale stable diffusion (SD) models [37], which are pre-trained on text-to-image (T2I) tasks, have demonstrated impressive semantic understanding and have been successfully applied to many downstream tasks [66, 35], including SR [46, 30, 59, 57, 56, 42]. While SD-based SR approaches have exhibited perceptually more realistic SR results than GAN-based methods [7], they commonly entangle the pixel-level fidelity and semantic-level enhancement objectives in the diffusion process [46, 57, 56], which can be contradictory in optimization [3, 43]. As a result, they struggle to balance the pixel-wise fidelity and semantic-level perception in the final SR output. While some methods [30, 59, 42] perform pixel-level restoration before semantic-level enhancement, the performance of the latter stage depends heavily on the accuracy of the earlier stage.
@@
Besides the trade-off between pixel-level fidelity and semantic -level details, users in practice often have different preferences on the SR results: some prioritize content fidelity over detail generation, while others favor semantically rich details over pixel-wise fidelity. This diversity in user preferences underscores the importance and demands for a more flexible SR approach that can accommodate individual tastes during inference. Although some previous methods [11, 34] have explored interactive SR methods to control the restoration strength, they are primarily limited to degradation-level adjustments (such as noise and blur), resulting in either smoother or sharper outputs. Several multi-step SD-based SR methods [57, 59, 30, 61] employ guidance-based strategies [14] at each sampling step to achieve varying levels of semantic richness, but these methods often struggle with precise control and efficiency.
@@
In this paper, we present Pixel-level and Semantic-level Adjustable Super-Resolution, namely PiSA-SR, which separates pixel-level and semantic-level enhancement into two distinct Low-Rank Adapter (LoRA) [16] weight spaces with the pre-trained SD model, providing an effective SR model for various user preferences. We begin by formulating the SD-based SR problem to learn the residual between LQ and HQ latent features [12, 64, 69]. With this formulation, we can not only accelerate the convergence of model training but also introduce scale factors on the model outputs, enabling flexible adjustment on the SR results during the inference stage without re-training. Then, we fine-tune the pre-trained SD model by introducing two separate LoRA modules dedicated to pixel-level regression and semantic-level enhancement. In specific, the $\ell_{2}$ loss is employed for the LoRA module of pixel-level regression, while the LPIPS [67] and classifier score distillation (CSD) [62] losses are employed for another LoRA module, leveraging the semantic priors encoded in the pre-trained VGG classification model [39] and SD image generation model [37]. A decoupled training approach is proposed to train pixel- and semantic-level LoRAs, effectively improving semantic information while maintaining pixel-level fidelity.
@@
Our experiments demonstrate that PiSA-SR not only achieves superior SR performance to existing SD-based models but also offers an effective way for users to adjust the SR styles based on their tastes. An example is shown in Fig. 1. The horizontal and vertical axes indicate the semantic-level and pixel-level guidance scales, respectively. Increasing the pixel-level scale removes noise and compression artifacts, but a too-strong scale over-smoothes image details. Conversely, increasing the semantic-level scale enriches the image details. Nonetheless, a too-high semantic scale over-enhances visual artifacts. PiSA-SR provides the flexibility to adjust both pixel-level and semantic-level guidance scales based on user preferences.
@@ 2 Related Work
Image super-resolution. Early deep learning-based SR methods [5, 10, 22, 27] aim to improve image fidelity metrics such as PSNR and SSIM [52]. By using bicubic downsampling to generate LQ-HQ image training pairs, various approaches have been developed to enhance the SR performance, such as dense [44], residual [69], recursive connections [22], non-local strategies [47], and attention mechanisms [27, 68, 8]. Though these advancements have significantly improved the SR performance, the real-world LQ images with complex degradations are much more challenging to enhance. While researchers have proposed to collect real-world LQ-HQ paired datasets [4, 55] to train SR models, simulating realistic LQ-HQ pairs is a more economic way [65, 50]. BSRGAN [65] randomly shuffles some basic degradation operators to synthesize LQ-HQ pairs, while RealESRGAN [50] implements high-order degradation modeling. Compared with pixel-level losses, the perceptual loss [17] and GAN loss [24, 48] are more effective in improving image perceptual quality. However, the adversarial training also leads to unnatural visual artifacts in the SR results. Many of the later works aim to reduce GAN-generated artifacts from the perspectives of frequency division [25], network optimization [70], the weighting of different losses [49, 43], and image local statistics [29].
@@
Diffusion model based super-resolution. Recent studies have explored the use of diffusion models (DMs) [15, 37] for SR. Early approaches adjust pre-trained DMs’ reverse transitions [20] using gradient descent, offering training-free solutions by assuming predefined image degradations. ResShift [63] trains a DM from scratch on paired LQ-HQ data. Recently, SD models pre-trained on T2I tasks have been widely used for SR tasks due to their strong image priors [46, 30, 59, 57, 56, 51, 32, 2, 36, 6]. StableSR [46] introduces a trainable encoder and takes LQ image as a condition in SD. DiffBIR [30] first applies a restoration module to mitigate degradations, then utilizes SD to enhance details. PASD [59] employs an encoder for degradation removal and introduces a pixel-aware cross-attention module to incorporate low-level and high-level image features into the SD process. SeeSR [57] enhances semantic robustness through degradation-aware tag-style prompts to guide diffusion. Considering that the multi-step diffusion process will increase the computational cost and the risk of synthesizing unfaithful contents, researchers started to develop fewer-step DM-based SR models. SinSR [51] applies consistency-preserving distillation to shorten the diffusion process of ResShift. OSEDiff [56] employs the LQ image as direct input, eliminating the random noise sampling. It adopts the VSD loss [54, 60] to distill the generative capacity from the multi-step diffusion process, offering a one-step DM-based SR solution. However, all the above-mentioned methods face a contradiction between pixel-level and semantic-level enhancement and most of them lack the flexibility to adjust the SR style for diverse user preferences.
@@ 3.1 Model Formulation
Learning the residual between the LQ and GT features has been successfully applied in deep learning-based restoration methods [64, 69]. However, DMs typically perform multi-step iterations, making applying residual learning difficult. The recently developed OSEDiff [56] employs a single diffusion step to achieve end-to-end SR training, making it feasible to adopt the residual learning strategy. Here, we formulate the SR problem as learning the residual between $z_{L}$ and $z_{H}$, as illustrated in Fig. 2 (c). We use subtraction in the global residual connection because DMs are trained to denoise in the inverse diffusion process. Such a residual learning formulation helps the model focus on learning the desired high-frequency information from the HQ latent features, avoiding extracting less relevant information from the LQ latent. It can also accelerate the convergence of the model training process [12]. Furthermore, we introduce a scaling factor $\lambda$ to adjust the residual $\epsilon_{\theta}$ added to the LQ latent, which can be expressed as:
@@
The $\lambda$ is fixed as $1$ during training. Users can adjust the output in the inference stage by using a smaller $\lambda$ (to preserve more original contents) or a higher $\lambda$ (to enhance details more aggressively) based on their preferences.
@@ 3.2 Dual-LoRA Training
Previous SD-based SR methods mostly entangle the pixel-level and semantic-level enhancement in the training process [46, 56], making the balance of content fidelity and perpetual quality difficult. Some works sequentially perform pixel-level and semantic-level enhancement using two separate networks [30]. However, errors in the earlier stage may propagate to the next stage, limiting overall performance. Furthermore, using two separate networks increases the computational burden and memory requirements, making the approach less efficient. It has been successfully applied in customized T2I tasks using distinct LoRA modules for various generation styles [38, 18, 41]. Inspired by this, we propose a decoupled training approach by utilizing two LoRA modules under the pre-trained SD model, targeting pixel-level and semantic-level enhancement for the SR task respectively. Our approach, namely Dual-LoRA, introduces only a small number of additional parameters during training. These LoRA parameters can also be merged into the pre-trained SD model during inference.
@@
Figure 4: The model outputs with pixel-wise and semantic-level losses for a given LQ image.
@@ 3.3 Pixel-level and Semantic-level Losses
where $c$ is the text prompt, $t$ is the current timestep, and $\sigma_{t}=\sqrt{1-{{\bar{\alpha}}_{t}}}$.
The SD model, parameterized by $\theta$, model score functions of $\nabla_{x_{t}}\log p_{\theta}(x_{t}\mid c)$ and $\nabla_{x_{t}}\log p_{\theta}(x_{t})$ using noise predictions [40]. By taking the expectation of the gradient in Eq. 4 over all possible $t$, we have the CSD [14] loss, which is originally used in 3D generation tasks to optimize the posterior probability of rendered images aligning their semantic content with text prompts.
@@
Note that the VSD loss [54, 60], which also aims to align the distributions of enhanced and natural images within the latent space, has been verified effective for SR by OSEDiff [56]. The gradient of VSD loss can be written as:
@@ 3.4 The Inference Process of PiSA-SR
Fig. 3(b) shows the inference process of PiSA-SR. In the default inference setting (represented by the green dash-dot line), only the PiSA-LoRA, which merges the pixel-level and semantic-level LoRA modules, is used to process the input together with the pre-trained SD model, achieving state-of-the-art SR performance in just one-step diffusion (see our experiments in Section 4.3). To achieve flexible SR with diverse user preferences, we extend the strategy in Eq. (3) by introducing a pair of pixel and semantic guidance factors, denoted by $\lambda_{pix}$ and $\lambda_{sem}$, to control the SR results:
@@
In Eq. (8), $\epsilon_{\theta_{pix}}(z_{L})$ is the output with only pixel-level LoRA, while $\epsilon_{\theta_{P\!i\!S\!A}}(z_{L})$ is the output with both pixel and semantic level enhancement. Inspired by CFG [14], the difference between them, i.e., $\epsilon_{\theta_{P\!i\!S\!A}}(z_{L})-\epsilon_{\theta_{pix}}(z_{L})$, can well separate pixel-level degradations from semantic details (see the red box in Fig. 4). Therefore, as illustrated in Fig. 3(b) (represented by the black dashed line), by adjusting $\lambda_{pix}$ and $\lambda_{sem}$, we can control the contributions of pixel-level LoRA and semantic-level LoRA, and consequently deliver the SR results with different styles, as shown in Fig. 1.
@@ 4.1 Experimental Settings
Training settings. We train PiSA-SR upon SD 2.1-base for the ×$\times 4$ SR task. The two trainable LoRA modules are applied to the weights of all convolution and MLP layers. Both modules are initialized using a Gaussian distribution with a rank of 4. Following SeeSR [57] and OSEDiff [56], we employ LSDIR [26] and the first 10K images from FFHQ [19] as the training data. We generate paired training data using the degradation pipeline from RealESRGAN [50]. The batch size is 16 and the training patch size is 512×512. The Adam optimizer [23] is used with a learning rate 5e-5. The pixel-level and semantic-level LoRA modules undergo 4K and 8.5K training iterations.
@@
Test datasets and evaluation metrics. Following previous arts [46, 57, 56], we evaluate the competing methods using synthetic and real-world test data. The synthetic dataset comprises 3000 images cropped from DIV2K [1] with size $512\times 512$, degraded using RealESRGAN degradation pipeline [50]. The real-world data are center-cropped from RealSR [4] and DrealSR [55] datasets with size $128\times 128$ for LQ images and $512\times 512$ for HQ images. PSNR and SSIM [52], computed on the Y channel in the YCbCr space, are used to measure the fidelity of SR results; LPIPS [67] and DISTS [9], computed in the RGB space, are used to assess the perceptual quality of SR results; FID [13] evaluates the distance of distributions between GT and restored images; NIQE [33], CLIPIQA [45], MUSIQ [21] and MANIQA [58] evaluate the image quality without reference images. Additionally, a user study is conducted to validate the effectiveness of the proposed adjustable SR in the supplementary materials.